OpenStudy (anonymous):

WILL GIVE FAN AND MEDAL PLZ HELP Evaluate. (5 • 8 – 12)(36 ÷ 9 • 4)

OpenStudy (anonymous):

I think its 448? Maybe

OpenStudy (anonymous):

Is the dot a multiplication symbol? If so, 8-12 = -4 5 x -4 = -20 9 x 4 = 36 36 / 36 = 1 -20 x 1 = -20.

OpenStudy (anonymous):

Yes, its a multiplication symbol, but am I correct that its 448?

OpenStudy (anonymous):

I think you are right, I did it and got the same answer because you have to follow PEMDAS and it would work out. Btw, LOVE your profile picture!!!! #10Rose

OpenStudy (anonymous):

Yay, I'm right for once! XD And thanks :) Haha, I love David Tennant X Rose Tyler, ship them so hard XD